1. Using the table below, find the standard deviation and the variance of the scores of 30 students in Statistics exam.





Class Interval Class Frequency (f)



47 - 51 4



52 – 56 3



57 – 61 3



62 – 66 4



67 – 71 5



72 – 76 3



77 – 81 3



82 – 86 1



87 - 91 4





2. Determine the standard deviation and the variance of the scores obtained by the students in Mathematics using the table below.





Class Interval Class Frequency (f)



29 – 33 5



34 – 38 5



39 – 43 4



44 – 48 4



49 – 53 2



54 – 58 4



59 – 63 6



64 - 68 2





1
Expert's answer
2022-01-16T12:54:38-0500

1.

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 & Classes & M_i & f_i & M_i\cdot f_i & M_i^2\cdot f_i \\ \hline
 & 47-51 & 49 & 4 & 196 & 9604 \\
 \hdashline
 & 52-56 & 54 & 3 & 162 & 8748 \\
 \hdashline
 & 57-61 & 49 & 3 & 177 & 10443 \\
 \hdashline
 & 62-66 & 64 & 4 & 256 & 16384 \\
 \hdashline
 & 67-71 & 69 & 5 & 345 & 23805 \\
 \hdashline
 & 72-76 & 74 & 3 & 222 & 16428 \\
 \hdashline
 & 77-81 & 79 & 3 & 237 & 18723 \\
 \hdashline
 & 82-86 & 84 & 1 & 84 & 7056 \\
 \hdashline
 & 87-91 & 89 & 4 & 356 & 31684 \\
 \hdashline
 Sum= & & & 30 & 2035 & 142875 \\
 \hdashline
\end{array}"


"\sigma^2=\dfrac{1}{N}\bigg(\sum _iM_i^2\cdot f_i-\dfrac{1}{N}\big(\sum_iM_i\cdot f_i\big)^2\bigg)"

"=\dfrac{1}{30}(142875-\dfrac{1}{30}(2035)^2)"

"\approx161.1389"

"\sigma=\sqrt{\sigma^2}=\sqrt{161.38888889}\approx12.6940"

2.

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 & Classes & M_i & f_i & M_i\cdot f_i & M_i^2\cdot f_i \\ \hline
 & 29-33 & 31 & 5 & 155 & 4805 \\
 \hdashline
 & 34-38 & 36 & 5 & 180 & 6480 \\
 \hdashline
 & 39-43 & 41 & 4 & 164 & 6724\\
 \hdashline
 &44-48 & 46 & 4 & 184 & 8464 \\
 \hdashline
 & 49-53 & 51 & 2 & 102 & 5202 \\
 \hdashline
 & 54-58 & 56 & 4 & 224 & 16428 \\
 \hdashline
 & 59-63 & 61 & 6 & 366 & 22326 \\
 \hdashline
 & 64-68 & 66 & 2 & 132 & 8712 \\
 \hdashline
 & 87-91 & 89 & 4 & 356 & 31684 \\
 \hdashline
 Sum= & & & 32 & 1507 & 75257 \\
 \hdashline
\end{array}"


"\sigma^2=\dfrac{1}{N}\bigg(\sum _iM_i^2\cdot f_i-\dfrac{1}{N}\big(\sum_iM_i\cdot f_i\big)^2\bigg)"

"=\dfrac{1}{32}(75257-\dfrac{1}{32}(1507)^2)"

"=133.9599609375\approx133.9600"

"\sigma=\sqrt{\sigma^2}=\sqrt{133.9599609375}\approx11.5741"